# A General Control Method for Human-Robot Integration

### v1

@@ Abstract
1 Introduction
Assistive and rehabilitation devices such as powered wheelchairs, assistive robotic arms, and limb prostheses play a crucial role in assisting individuals with severe motor impairments Gopinath et al. (2017), which require daily assistance due to e.g. spinal cord or brain injuries.
The chosen interfaces, such as joysticks, head arrays, and sip-and-puff, often pose challenges in managing assistive devices Jain et al. (2015). They have an impact on aspects such as timing, accuracy, and transient noise Edakkattil Gopinath et al. (2020). These interfaces are low dimensional and operate discretely, necessitating users to switch between different areas of the control space. Partially autonomous control can reduce cognitive and physical burden on users Gopinath et al. (2017).
@@ 1 Introduction
In addition to techniques that exploit Inertial Measurement Unit (IMU) sensors in Lee et al. (2021) and brain-machine interfaces (BMIs) in Hochberg et al. (2012) to control assistive devices with high number of Degrees of Freedom (DoFs), another approach involves employing dimensionality reduction techniques, such as Principal Component Analysis (PCA), Hochberg et al. (2012). Beyond the branch of assistive robotics, these control interfaces find applications also in human-robot interfaces (HRIs), for both domestic and professional settings. Several studies implement different interfaces to control robotic devices, such as upper-body movements Sanna et al. (2013), or electromyographic sensors (EMG) on the forearm Stoica et al. (2014), or sensorised upper-body soft exoskeleton, coupled with virtual reality goggles, Stoica et al. (2014), or joint-level force control as the ones for the supernumerary robotic limbs Prattichizzo et al. (2021). Another interface system, which includes also a haptic admittance module, is presented in Lamon et al. (2020).
In the prosthetic field, the use of EMG sensors is the most widely adopted technique Farr et al. (2018). One of the main drawbacks is that as the level of amputation rises, a greater number of sensors is required to provide more inputs, while less independent measurement sites are available. Additionally, a strategy becomes necessary to switch the controller between multiple DoFs. Furthermore, the user’s skin condition can impact control performance, as can changes in electrode positions Roche et al. (2014). This control complexity may lead prosthetic users to develop undesired compensatory motion habits, which in turn, can result in postural issues and chronic pain Carey et al. (2008).
In recent years, researchers considered that compensatory motions of the patient’s body could be themselves useful to control prostheses. Maimeri et al. (2019) used IMUs on the stump in combination with surface EMG signals to control multi-channel prostheses and supernumerary limbs. Legrand et al. (2020); Merad et al. (2020); Legrand et al. (2022) more explicitly introduced the idea of exploiting body compensation as a signal to control upper limb prostheses.
@@
The generalization of the framework allows us to consistently include widely varying systems, both in the human component, where different abilities can be accommodated for, and in the artificial part, where upper-limb prostheses can be extended to assistive devices and even remote whole-body avatars. Preliminary results in this direction were presented in Feder et al. (2024). In this paper, we substantially extend and improve the framework introducing a model of the human reaching and compensation control, and an observer which can predict the reaching target intention and correct the controlled motion on–line.
@@ 2 Problem Definition
The task for us is to control $\dot{q}_{r}$ so that $e_{e}$ tends to zero, and to exploit a possible redundancy in the kinematics of the human-robot system to favor a relaxed configuration of the human joints $q_{h}$. The problem is not trivial because the desired final position $\bar{x}_{e}$ is not available to the robot controller. To circumvent this problem, we exploit the ability of humans to include in their body schema devices and appendages Merleau-Ponty (2010), and to exploit them to reach their goals. For example, a person wearing even a completely passive prosthesis, such as a “cosmetic hand”, can easily reach for an object, although at the cost of possibly exaggerated motions of the intact body joints in the shoulder, torso, waist and legs.
Consider hence the case that the user attempts to reach the goal without counting on the assistance of an active prosthesis, as they would do with a passive appendage of some form: the user will employ their own DoFs ($q_{h}$) while assuming that the prosthetic joints are fixed ($\dot{q}_{r}=0$) in the present configuration. The resulting body motions (which are somehow not natural and possibly fatiguing for the user) are commonly referred to as compensatory movements. The idea of exploiting compensatory motions in prosthesis control was first introduced by Legrand et al. (2022). Let $x_{c}$ denote a compensatory reference frame fixed on a selected body segment (e.g. at the shoulder), let $\bar{x}_{c}$ indicate its posture in a relaxed body posture, and define a compensation error $e_{c}=\bar{x}_{c}-x_{c}$.
@@ 3 Approach
3.1 Control of Physically Connected Human–Robot Systems
@@
3.2 Control of Physically Disconnected Human–Robot Systems
@@
Indeed, because of the physical robot-user disconnection, in (6) it holds $J_{he}=0$. Should also the human model in (7) include $\hat{J}_{he}=J_{he}=0$, the operator could only try to solve the compensatory task, and not act on the robot at all. However, this contrasts with common experience in observing pilots in virtual reality and in tele-operation systems Betancourt et al. (2023), and even in commercial videogames. Indeed human operators often move their bodies and try (sometimes even very hard) to move the end-effector toward a task as if they were the avatar to which the end-effector belongs. In other words, the pilot immerses into the remote environment through the perspective of the robot avatar, usually by viewing the environment through the robot’s eyes (e.g. via a head mounted display (HMD) system). This phenomenon could be regarded as a generalized version of the rubber hand effect, a rather well studied illusion in the psycological literature (see e.g. Flögel et al. (2014)).
@@ 4 Numeric Analysis
4.1 Simulations of Physically Connected Human-Robot System
@@
4.2 Simulation of Physically Disconnected Human-Robot System
@@
For this study, we employed the Alter-Ego robot Zambella et al. (2019), a two-wheeled auto-balancing humanoid avatar. The robot has a mobile base for translation and rotation, along with two robotic arms, each with 5 DoFs, attached to the main robotic body. Considering its kinematics, we designated the matrix $J_{re}$ as the Jacobian matrix of the entire robotic chain from the global frame (fixed to the ground) to the right end effector, thus being influenced by the joints of the wheels and the right arm. The robotic base is described by the $x$ and $y$ coordinates of the center of its base, and by the orientation angle $\theta$ around the vertical axis $z$. The base motions are modeled as a unicycle, with the translation and rotation velocities as inputs.
@@ 5 Experiments and Results
5.1 Experiments with a Virtual Upper-Limb Prosthesis
@@
5.2 Experiment with a Whole-Body Prosthesis
In the second experiment, we validate the algorithm for the two-wheeled humanoid robotic avatar Alter-Ego Zambella et al. (2019). Robotic joints to be controlled here include not only those belonging to the arm (5 joints for the right arm) but also the wheeled base. The base is modelled by the kinematics of a unicycle, as described in Section 4.
@@ 5.2 Experiment with a Whole-Body Prosthesis
6 Discussion
Our research moved from previous work Legrand et al. (2020) showing how users of simple prosthetic devices tend to make up for missing DoFs by performing so-called compensatory motions with the functional parts of their body, and how these motions can be used to drive the active prosthetic limbs and reduce compensation.
@@ 6 Discussion
Possible applications of the methods described here are to assist people with disabilities that limit their motions, such as those with neurological disorders. Preliminary tests (not described here) showed that a potential issue in this case are involuntary reflex motions these individuals might experience. Implementing an algorithm for motion discrimination can significantly help in this context. The implementation would distinguish between intentional movements and involuntary reflexes, ensuring that only purposeful motions trigger the robotic commands. Additionally, integrating a deadband zone around the user’s target posture can facilitate controller deactivation, particularly when significant residual motions are present. This deadband zone can also help in discriminating voluntary motions from the involuntary ones of the shoulders, associated from example with breathing Farshchiansadegh et al. (2014). Further research into incorporating kinematics studies in the control process could improve robotic movements, enhancing overall system performance.

### v2

@@ Abstract
I Introduction
Assistive and rehabilitation devices such as powered wheelchairs, assistive robotic arms, and limb prostheses play a crucial role in assisting individuals with severe motor impairments [1], which require daily assistance due to e.g. spinal cord or brain injuries.
The chosen interfaces, such as joysticks, head arrays, and sip-and-puff, often pose challenges in managing assistive devices [2]. They have an impact on aspects such as timing, accuracy, and transient noise [3]. These interfaces are low dimensional and operate discretely, necessitating users to switch between different areas of the control space. Partially autonomous control can reduce cognitive and physical burden on users [1].
@@ I Introduction
In addition to techniques that exploit Inertial Measurement Unit (IMU) sensors in [4] and brain-machine interfaces (BMIs) in [5] to control assistive devices with high number of Degrees of Freedom (DoFs), another approach involves employing dimensionality reduction techniques, such as Principal Component Analysis (PCA), [5]. Beyond the branch of assistive robotics, these control interfaces find applications also in human-robot interfaces (HRIs), for both domestic and professional settings. Several studies implement different interfaces to control robotic devices, such as upper-body movements [6], or electromyographic sensors (EMG) on the forearm [7], or sensorised upper-body soft exoskeleton, coupled with virtual reality goggles, [7], or joint-level force control as the ones for the supernumerary robotic limbs [8]. Another interface system, which includes also a haptic admittance module, is presented in [9].
In the prosthetic field, the use of EMG sensors is the most widely adopted technique [10]. One of the main drawbacks is that as the level of amputation rises, a greater number of sensors is required to provide more inputs, while less independent measurement sites are available. Additionally, a strategy becomes necessary to switch the controller between multiple DoFs. Furthermore, the user’s skin condition can impact control performance, as can changes in electrode positions [11]. This control complexity may lead prosthetic users to develop undesired compensatory motion habits, which in turn, can result in postural issues and chronic pain [12].
In recent years, researchers considered that compensatory motions of the patient’s body could be themselves useful to control prostheses. [13] used IMUs on the stump in combination with surface EMG signals to control multi-channel prostheses and supernumerary limbs. [14, 15, 16] more explicitly introduced the idea of exploiting body compensation as a signal to control upper limb prostheses.
@@
The generalization of the framework allows us to consistently include widely varying systems, both in the human component, where different abilities can be accommodated for, and in the artificial part, where upper-limb prostheses can be extended to assistive devices and even remote whole-body avatars. Preliminary results in this direction were presented in [17]. In this paper, we substantially extend and improve the framework introducing a model of the human reaching and compensation control, and an observer which can predict the reaching target intention and correct the controlled motion on–line.
@@ II Problem Definition
The task for us is to control $\dot{q}_{r}$ so that $e_{e}$ tends to zero, and to exploit a possible redundancy in the kinematics of the human-robot system to favor a relaxed configuration of the human joints $q_{h}$. The problem is not trivial because the desired final position $\bar{x}_{e}$ is not available to the robot controller. To circumvent this problem, we exploit the ability of humans to include in their body schema devices and appendages [18], and to exploit them to reach their goals. For example, a person wearing even a completely passive prosthesis, such as a “cosmetic hand”, can easily reach for an object, although at the cost of possibly exaggerated motions of the intact body joints in the shoulder, torso, waist and legs.
Consider hence the case that the user attempts to reach the goal without counting on the assistance of an active prosthesis, as they would do with a passive appendage of some form: the user will employ their own DoFs ($q_{h}$) while assuming that the prosthetic joints are fixed ($\dot{q}_{r}=0$) in the present configuration. The resulting body motions (which are somehow not natural and possibly fatiguing for the user) are commonly referred to as compensatory movements. The idea of exploiting compensatory motions in prosthesis control was first introduced by [16]. Let $x_{c}$ denote a compensatory reference frame fixed on a selected body segment (e.g. at the shoulder), let $\bar{x}_{c}$ indicate its posture in a relaxed body posture, and define a compensation error $e_{c}=\bar{x}_{c}-x_{c}$.
@@ III Approach
III-A Control of Physically Connected Human–Robot Systems
@@
III-B Control of Physically Disconnected Human–Robot Systems
@@
Indeed, because of the physical robot-user disconnection, in (6) it holds $J_{he}=0$. Should also the human model in (7) include $\hat{J}_{he}=J_{he}=0$, the operator could only try to solve the compensatory task, and not act on the robot at all. However, this contrasts with common experience in observing pilots in virtual reality and in tele-operation systems [19], and even in commercial videogames. Indeed human operators often move their bodies and try (sometimes even very hard) to move the end-effector toward a task as if they were the avatar to which the end-effector belongs. In other words, the pilot immerses into the remote environment through the perspective of the robot avatar, usually by viewing the environment through the robot’s eyes (e.g. via a head mounted display (HMD) system). This phenomenon could be regarded as a generalized version of the rubber hand effect, a rather well studied illusion in the psycological literature (see e.g. [20]).
@@ IV Numeric Analysis
IV-A Simulations of Physically Connected Human-Robot System
@@
IV-B Simulation of Physically Disconnected Human-Robot System
@@
For this study, we employed the Alter-Ego robot [21], a two-wheeled auto-balancing humanoid avatar. The robot has a mobile base for translation and rotation, along with two robotic arms, each with 5 DoFs, attached to the main robotic body. Considering its kinematics, we designated the matrix $J_{re}$ as the Jacobian matrix of the entire robotic chain from the global frame (fixed to the ground) to the right end effector, thus being influenced by the joints of the wheels and the right arm. The robotic base is described by the $x$ and $y$ coordinates of the center of its base, and by the orientation angle $\theta$ around the vertical axis $z$. The base motions are modeled as a unicycle, with the translation and rotation velocities as inputs.
@@ V Experiments and Results
V-A Experiments with a Virtual Upper-Limb Prosthesis
@@
V-B Experiment with a Whole-Body Prosthesis
In the second experiment, we validate the algorithm for the two-wheeled humanoid robotic avatar Alter-Ego [21]. Robotic joints to be controlled here include not only those belonging to the arm (5 joints for the right arm) but also the wheeled base. The base is modelled by the kinematics of a unicycle, as described in Section IV.
@@ V-B Experiment with a Whole-Body Prosthesis
VI Discussion
Our research moved from previous work [14] showing how users of simple prosthetic devices tend to make up for missing DoFs by performing so-called compensatory motions with the functional parts of their body, and how these motions can be used to drive the active prosthetic limbs and reduce compensation.
@@ VI Discussion
Possible applications of the methods described here are to assist people with disabilities that limit their motions, such as those with neurological disorders. Preliminary tests (not described here) showed that a potential issue in this case are involuntary reflex motions these individuals might experience. Implementing an algorithm for motion discrimination can significantly help in this context. The implementation would distinguish between intentional movements and involuntary reflexes, ensuring that only purposeful motions trigger the robotic commands. Additionally, integrating a deadband zone around the user’s target posture can facilitate controller deactivation, particularly when significant residual motions are present. This deadband zone can also help in discriminating voluntary motions from the involuntary ones of the shoulders, associated from example with breathing [22]. Further research into incorporating kinematics studies in the control process could improve robotic movements, enhancing overall system performance.